Consider the titration of 100.0 mL of 0.100 M /hydrazine () by 0.200 M . Assume that hydrazine is monoprotic. Determine the following: a. pH before any HNO3 is added b. volume of acid to reach the equivalence point c. pH when 20.0 mL of HNO3 has been added d. pH when 25.0 mL of HNO3 has been added e. pH when 40.0 mL of HNO3 has been added f. pH when 50.0 mL of HNO3 has been added g. pH when 100.0 mL of HNO3 has been added h. Draw a titration curve for this titration and indicate on the curve points a, c, d, e, f, and g i. Draw a titration curve for this titration but assuming that hydrazine is a strong base. You do not have to recalculate the values from a-f. Highlight on this curve the differences from the curve you drew for h. j. Assume instead that hydrazine can accept two protons. Which if your values from a-g would be different? For the values that are different, what are their new values?

***Really just need help with J***
Answer #1

[OH-] H 2020-22 mec rsumed 02-40 mel 01 60x0L. 120 OOS 92 Llytherine] = 0%(%) -- [uri) = n = 109%) (pH--600) = 13) comes figSUU by Volume of Paid to Reach He Equivalence Points 10020 1 - Sont of HINO di pH=7 at Equivalence Point, because [Hydrazine]halo g Strong Base : Weak Base 0 Weak Base Vol. of H10,(htaAN DO ogs Luv) comb 10 = 13:29 by volume of HNO3 Required = 1100m Oy Hyobrazina 7= 5020T = 0-067632. [1] . 10-19 9 Hydrazineox[Hydrazine). I = 0486). 18Hz 12-93 By [Hydrazine) - S0xOcl30-032m).DH=112182 ga bit=7 at Equivalence Point